OpenStudy (anonymous):

can anyone help me with this? The density of an aqueous solution containing 12.50 g K2SO4 in 100.00 g solution is 1.083 g/mL. Calculate the concentration of this solution in terms of molarity, molality, percent of K2SO4, and mole fraction of solvent. i don't know how to get the liters of solution

OpenStudy (anonymous):

solute-K2SO4 My solution on Molarity is M=moles of solute/liters of solution moles of K2SO4= 12.50g/174.2592g/mol = 0.0717 mol. M= 0.0717mol/ Liters of solution

OpenStudy (joannablackwelder):

100 g solution ( 1 mL/1.083 g) = ? mL solution?

OpenStudy (joannablackwelder):

Density = mass/volume so Volume = mass/density

OpenStudy (anonymous):

You've got 100.00 g of solution with density 1.083 g/mL, so the volume of solution is: (100.00 g) / (1.083 g/mL) = 92.336 mL (12.50 g K2SO4) / ( 174.2592 g K2SO4/mol) / (0.092336 L) = 0.7769 mol/L K2SO4 (12.50 g K2SO4) / ( 174.2592 g K2SO4/mol) / (100.00 g) x (1000 g) = 0.7173 m (12.50 g K2SO4) / (100.00 g) = 12.50% K2SO4 by mass (100.00 g total) - (12.50 g K2SO4) = 87.50 g H2O (87.50 g H2O) / (18.01532 g H2O/mol) = 4.85698 mol H2O (12.50 g K2SO4) / ( 174.2592 g K2SO4/mol) = 0.0717322 mol K2SO4 (4.85698 mol) / (4.85698 mol + 0.0717322 mol) = 0.9854 mol fraction of H2O

OpenStudy (anonymous):

thank you :)